OpenStudy (anonymous):

verify the identity 1 + (sec^2)x *(sin^2)x = (sec^2)x

OpenStudy (anonymous):

what's the most fundamental trig identity

OpenStudy (anonymous):

cos^2x+sin^2x=1 so (if you divide both sides by cos^2x)... 1+tan^2x=sec^2x tan^2x=sec^2x-1 sec^2x/sec^2x-1=csc^2x sec^2x/tan^2x=csc^2x (1/cos^2x)/(sin^2x/cos^2x)=csc^2x (1/cos^2x)(cos^2x/sin^2x)=csc^2x (1/sin^2x)=csc^2x csc^2x=csc^2x

OpenStudy (anonymous):

Why is \[ (sec^2)x *(sin^2)x=\tan^2(x) \]

OpenStudy (anonymous):

*hint* most people are skipping this part, but i think it is crucial: sec(x) = 1/cos(x) sec^2(x) = 1/cos^2(x)

OpenStudy (anonymous):

So \[ 1+ \tan^2(x)=\sec^2(x) \]

OpenStudy (anonymous):

All trig functions are of x, deleted for brevity. Given cos^2 + sin^2 =1, multiply both sides by sec^2. Your conclusion follows directly.

OpenStudy (anonymous):

everyone's saying different things :/ im confused...

OpenStudy (anonymous):

Not really. Everything starts from sin^2 + cos^2 = 1

OpenStudy (anonymous):

Now it's true, you have to learn to forget about x or theta what angle the sine is for. And you also learn to forget about the radius, because this is unit circle.

OpenStudy (anonymous):

\[ 1 + \sec^2x \, \sin^2x =\\ 1 + \frac 1{\cos^2 x} \sin^2x= 1+\tan^2(x)=\sec^2 x \]

OpenStudy (anonymous):

Good work with the equation editor @eliassaab

OpenStudy (anonymous):

thank u everyone! :)

OpenStudy (anonymous):

That is a summary of what was posted above \[ 1 + \sec^2x \, \sin^2x =\\ 1 + \frac 1{\cos^2 x} \sin^2x= \frac {\sin^2 x+\cos^2 x}{\cos^2 x}= \frac 1{\cos^2(x)}=\sec^2(x) \]

OpenStudy (anonymous):

wait @ eliassaab i dont understand how you got sin2x+cos2x/cos2x

OpenStudy (anonymous):

same denominator

OpenStudy (anonymous):

of what?

OpenStudy (anonymous):

Here are more details \[ 1 + \sec^2x \, \sin^2x =\\ 1 + \frac 1{\cos^2 x} \sin^2x= \frac{ \cos^2 x}{\cos^2 x}+\frac{\sin^2x}{\cos^2 x}=\\ \frac {\sin^2 x+\cos^2 x}{\cos^2 x}= \frac 1{\cos^2(x)}=\sec^2(x) \]

OpenStudy (anonymous):

Did you get it now?

OpenStudy (anonymous):

where did the one go?

OpenStudy (anonymous):

\[ 1= \frac{ \cos^2 x}{\cos^2 x} \]

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

yes so (1/cos2x)*sin2x=sin2x/cos2x??

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@telliott99

OpenStudy (anonymous):

Correct, sarah.

OpenStudy (anonymous):

Sorry I was gone, but yes you are correct. Go back through and look carefully at @eliassaab stuff.

OpenStudy (anonymous):

ok. thank u everyone